### v1

# GPU-Accelerated Parallel Selected Inversion for Structured Matrices Using sTiles

Abstract
Selected inversion is essential for applications such as Bayesian inference, electronic structure calculations, and inverse covariance estimation, where computing only specific elements of large sparse matrix inverses significantly reduces computational and memory overhead. We present an efficient implementation of a two-phase parallel algorithm for computing selected elements of the inverse of a sparse symmetric matrix $A$, which can be expressed as $A=LL^{T}$ through sparse Cholesky factorization. Our approach leverages a tile-based structure, focusing on selected dense tiles to optimize computational efficiency and parallelism. While the focus is on arrowhead matrices, the method can be extended to handle general structured matrices. Performance evaluations on a dual-socket 26-core Intel Xeon CPU server demonstrate that sTiles [^1] https://github.com/esmail-abdulfattah/sTiles outperforms state-of-the-art direct solvers such as Panua-PARDISO, achieving up to 13X speedup on large-scale structured matrices. Additionally, our GPU implementation using an NVIDIA A100 GPU demonstrates substantial acceleration over its CPU counterpart, achieving up to 5X speedup for large, high-bandwidth matrices with high computational intensity. These results underscore the robustness and versatility of sTiles, validating its effectiveness across various densities and problem configurations.
Index Terms:
Sparse Matrix Computations, Arrowhead Structured Matrices, Tile Algorithms, Incomplete Inverse, Partial Inversion.
I Introduction
Matrix inversion is a fundamental operation in numerical linear algebra, which is pivotal to numerous applications in science and engineering. However, inverting large, sparse symmetric matrices is computationally intensive, particularly when dealing with specialized structures like arrowhead matrices. These matrices, characterized by non-zero elements concentrated along the block diagonal, the last block row, and the last block column, are prevalent in fields such as mathematics, physics, and engineering.
Traditional inversion methods often compute the entire inverse matrix, leading to a loss of sparsity as originally sparse structures are transformed into dense ones. This phenomenon can be interpreted through the lens of the Cayley-Hamilton theorem, which states that every square matrix satisfies its characteristic equation. As a result, when inverting matrices symbolically or computationally, terms in the matrix power series contribute to non-zero elements in all positions, resulting in the proliferation of fill-ins. Consequently, this highlights the inefficiency of direct inversion approaches for sparse matrices and motivates the need for tailored techniques that preserve sparsity.
Figure 1: Top row: matrix A (left) and the selected inverse of matrix A (right). Bottom row: matrix B (left) and the selected inverse of matrix B (right).
To address this challenge, selected inversion techniques have emerged as a focused alternative, computing only specific entries or substructures of the inverse matrix that are required for the target application. By avoiding the computation of elements considered unnecessary, the selected inversion significantly reduces resource consumption. This makes it particularly valuable in domains such as electronic structure calculations [1], large-scale sparse inverse covariance estimation [2], and Bayesian modeling [3]. However, despite these advantages, achieving scalability in selected inversion methods remains a significant challenge, particularly for high-dimensional applications.
The sparsity patterns of arrowhead matrices offer unique computational advantages, particularly in Bayesian inference, where they model interactions between multiple random effects. These matrices are commonly constructed using Kronecker products, for example, by capturing space-time dependencies in spatio-temporal models. Figure 1 illustrates this by comparing two arrowhead matrices (matrix A and matrix B) with differing sparsity levels. The sparsity level of these matrices often reflects the number of spatial locations or time points used in a model. As data collection technologies advance, we are entering an era of increasingly high-resolution datasets, where more sensors, satellites, and monitoring systems generate vast amounts of spatial and temporal information, leading to lower levels of sparsity in arrowhead matrices, driven by higher resolution and wider matrix bandwidth, as the number of locations increases.
Arrowhead matrices of this form are widely used in various scientific domains, particularly in Bayesian modeling [4], including geosciences, where they are employed to model climate variations and assess geological risks [5]. In epidemiology, they play a crucial role in tracking the spread of infectious diseases over time and space [6], and have been adopted in public health efforts such as excess mortality estimation by the World Health Organization (WHO) [19] and analyses of teen birth rates and drug poisoning mortality by the Centers for Disease Control and Prevention (CDC) [20]. Biostatistics benefits from these matrices in patient health modeling, particularly for analyzing longitudinally and spatially distributed data [7]. They are also integral to environmental modeling (e.g., predicting air pollution levels and ecological trends [8] and global demographic studies such as the United Nations’ World Population Prospects 2024 [21]). Additionally, meteorology relies on them to model weather patterns based on historical spatio-temporal data.
Despite their structured sparsity, computing the full inverse of such matrices results in a fully dense matrix, leading to significant memory and computational overhead. However, by selectively computing only the inverse of the non-zero elements in matrices A and B, we preserve their arrowhead structure, reducing storage and computational costs. As illustrated in Figure 1, regardless of the initial sparsity level, the selected inverse retains a dense arrowhead shape. This highlights the need of working with fine-grained data structures (e.g., tiles) from the outset, as it enables optimized computation and memory management without unnecessarily transforming the problem into a fully dense form.
To our knowledge, there is limited literature on parallel selected inversion methods specifically designed for sparse symmetric matrices with an arrowhead structure that utilize tile-based computations and the $LL^{T}$ Cholesky factorization. Existing methods, such as the Supernodal Selected Inversion (SINV) algorithm [9], compute selected inverse elements by leveraging supernodal partitions from LU factorizations and traversing elimination trees in parallel. Similarly, the SelInv algorithm [10] targets sparse symmetric matrices and employs $LDL^{T}$ factorizations with block algorithms and supernodes for efficiency. While both approaches benefit from multithreaded BLAS-level parallelism, they face challenges with irregular memory access patterns and dependency management inherent to supernodes.
Additional advancements include the selected inversion method for block tridiagonal arrowhead (BTA) matrices introduced in [11], which builds on strategies from quantum transport simulations and Kalman-Bucy filtering to efficiently compute block diagonal elements. However, its reliance on the specific BTA structure limits its general applicability and parallelization potential. Likewise, the hybrid parallel approach in [3] combines Krylov subspace methods and domain decomposition to approximate selected inverse elements in spatio-temporal Gaussian Markov Random Fields (GMRFs) [12]. Although scalable, this method relies on approximations through sampling and heuristic parameters.
Parallel implementations of Takahashi-based selected inversion for general sparse matrices have been explored, including a distributed-memory approach by [13], which integrates block-based Schur complement decomposition and relies on sparse direct solvers (e.g., PARDISO [18]) for symbolic and numeric factorization. In contrast, our method targets structured sparse symmetric matrices (e.g., arrowhead form), and achieves exact parallel solutions through tile-based computations, static scheduling, and GPU acceleration.
Building on this foundation, tile methods provide finer granularity, enhanced data locality, and superior scalability, making them especially effective for structured matrices like arrowheads. sTiles [14], a tile-based framework for high-performance linear algebra, exemplifies these advantages by utilizing sparse-dense tile computations for efficient factorizations and subsequent matrix operations. By design, sTiles minimizes inter-task dependencies and optimizes parallel execution, forming a robust foundation for extending selected inversion techniques. In this work, we extend the functionality of sTiles to efficiently perform selected inversion, significantly broadening its practical applicability to broader matrix structures.
This article is organized as follows. Section 2 introduces a recursive tile-based inversion algorithm built upon Cholesky decomposition. In Section 3, we detail a two-phase parallel algorithm, highlighting its static scheduling approach and GPU implementation with a focus on arrowhead structures. Section 4 presents performance results, including comparisons with existing libraries and the effectiveness of GPU implementation. Finally, we conclude with a summary of our findings and discuss potential future directions.
II Recursive Tile-Based Inversion via Cholesky Decomposition
Efficient inversion of large symmetric positive-definite matrices is central to many scientific applications, particularly when only selected entries of the inverse are needed. By leveraging the Cholesky factorization and operating at the tile level, we enable an approach that exploits data locality, facilitates parallelism, and scales well across hardware platforms.
Cholesky decomposition provides a natural foundation for this strategy. Given a matrix A, the decomposition
$A=LL^{T},$
where $L$ is a lower triangular matrix, forms the basis for many matrix inversion algorithms. Traditional approaches compute the inverse of $A$ by inverting $L$ elementwise and then computing $(L^{-1})^{T}$, but these methods are often inefficient for large matrices. To overcome these limitations, tile-based algorithms decompose $L$ into smaller blocks or tiles, enabling parallel computation of the inverse while managing interdependencies between tiles to maintain accuracy.
Early work by Agullo et al. introduced a tile-based in-place algorithm for the full inversion of symmetric positive-definite matrices, leveraging dynamic scheduling and compiler-inspired techniques such as loop reversal and array renaming to improve parallelism on multicore architectures [15]. This laid the foundation for asynchronous, task-based inversion strategies in dense linear algebra.
To illustrate, the derivation of the full inversion is based on the relationship
$L^{T}\Sigma=L^{-1},$
where both $L$ and $\Sigma$ are represented in a tile-based structure. For simplicity, consider $3\times 3$ tiles. The transpose of the matrix $L$ is given as:
$L^{T}=\begin{bmatrix}L_{00}&L_{10}&L_{20}\\ 0&L_{11}&L_{21}\\ 0&0&L_{22}\end{bmatrix},\quad\Sigma=\begin{bmatrix}\Sigma_{00}&\Sigma_{01}&\Sigma_{02}\\ \Sigma_{10}&\Sigma_{11}&\Sigma_{12}\\ \Sigma_{20}&\Sigma_{21}&\Sigma_{22}\end{bmatrix},$
where $\Sigma_{ij}$ and $L_{ij}$ are the submatrices corresponding to the $(i,j)$-th tile in the respective matrices. Each tile can be either dense or sparse, depending on the matrix structure and sparsity pattern.
We start the recursive computation with the tile $\Sigma_{22}$, using the relationship:
$L_{22}^{T}\Sigma_{22}=L_{22}^{-1},\quad\Sigma_{22}=L_{22}^{-T}L_{22}^{-1}.$
Next, we compute $\Sigma_{12}$ using:
$\Sigma_{21}L_{11}+\Sigma_{22}L_{21}^{T}=0,\quad\Sigma_{21}=-\Sigma_{22}L_{21}^{T}L_{11}^{-1}.$
These computations propagate recursively to compute other tiles of $\Sigma$. The relationships for each tile are summarized below:
$\begin{array}[]{rl}\Sigma_{22}&=L_{22}^{-T}L_{22}^{-1},\\[6.0pt] \Sigma_{21}&=-\Sigma_{22}L_{21}^{T}L_{11}^{-1},\\[6.0pt] \Sigma_{11}&=-\Sigma_{12}L_{21}^{T}L_{11}^{-1}+L_{11}^{-T}L_{11}^{-1},\\[6.0pt] \Sigma_{20}&=-\Sigma_{21}L_{10}^{T}L_{00}^{-1}-\Sigma_{22}L_{20}^{T}L_{00}^{-1},\\[6.0pt] \Sigma_{10}&=-\Sigma_{11}L_{10}^{T}L_{00}^{-1}-\Sigma_{12}L_{20}^{T}L_{00}^{-1},\\[6.0pt] \Sigma_{00}&=-\Sigma_{01}L_{10}^{T}L_{00}^{-1}-\Sigma_{02}L_{20}^{T}L_{00}^{-1}+L_{00}^{-T}L_{00}^{-1}.\end{array}$
The equations demonstrate how the inverse tiles are computed recursively, starting from the bottom-right corner of $\Sigma$ and propagating through the dependencies to the top-left corner. Algorithm 1 outlines the tile-based inversion procedure for a full matrix $A$ using its Cholesky decomposition $A=LL^{T}$.
The algorithm starts by computing diagonal tiles $\Sigma_{ii}$ using the relationship $\Sigma_{ii}=L_{ii}^{-T}L_{ii}^{-1}$, followed by updating off-diagonal tiles $\Sigma_{ji}$ through recursive propagation of dependencies. The key operations involve matrix multiplications and additions performed at the tile level, ensuring efficient computation.
Algorithm 1 Tile-based inversion of a full matrix
Initialization:
int i, j, k;
for $i=N-1$ to 00 step $-1$ do
     for $j=N-1$ to $i$ step $-1$ do
         if $i==j$ then
              $\Sigma_{ii}\leftarrow\Sigma_{ii}+L_{ii}^{-T}L_{ii}^{-1}$
              for $k=i+1$ to $N$ do
                  $\Sigma_{ii}\leftarrow\Sigma_{ii}-\Sigma_{ik}L_{ki}L_{ii}^{-1}$
              end for
         else
              $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jj}L_{ji}L_{ii}^{-1}$
              for $k=i+1$ to $N$ do
                  if $j!=k$ then
                       $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jk}L_{ki}L_{ii}^{-1}$
                  end if
              end for
         end if
     end for
end for
The core tile operations required for recursive inversion are implemented using well-defined linear algebra kernels. These operations are as follows:
  • TRSM (Triangular Solve with Multiple Right-Hand Sides): This kernel computes the inverse of a diagonal tile by solving a triangular system with the identity matrix:
    $\Sigma_{ii}\leftarrow\Sigma_{ii}^{-1}.$
  • LAUUM (Lower Triangular Matrix Multiplication): This kernel updates the diagonal tile by computing the product of a lower triangular matrix and its transpose. The result is stored in the upper triangular part and then mirrored to the full tile:
    $\Sigma_{ii}\leftarrow\Sigma_{ii}\Sigma_{ii}^{T}.$
  • GEMM (General Matrix-Matrix Multiplication): This operation updates an off-diagonal tile by performing matrix multiplication and subtracting the result:
    $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{kj}^{T}\Sigma_{ki}.$
  • TRMM (Triangular Matrix-Matrix Multiplication): This kernel updates off-diagonal tiles by multiplying them with a triangular matrix:
    $\Sigma_{ji}\leftarrow L_{jj}\Sigma_{ji}.$
In this work, we build on the tile paradigm but focus on structured sparse matrices, extending the tile-based framework to selected inversion, where only specific entries of the inverse are computed. This selective approach introduces significant computational savings by avoiding unnecessary operations: if a given operation (e.g., computing a particular tile update) does not contribute to any of the user-requested tiles, it is skipped entirely. This stands in contrast to the full inversion approach outlined in Algorithm 1, where all tiles are processed regardless of necessity. In our implementation, we adapt this algorithm by incorporating a filtering mechanism that prunes irrelevant computations while preserving correctness.
This targeted computation is particularly beneficial for structured matrices, such as arrowhead or banded forms, where sparsity patterns can be exploited to further reduce overhead. Our approach adopts a static scheduling strategy and targets hybrid CPU/GPU systems for optimal performance. We also generalize the tile-based inversion framework introduced by [11] to support broader classes of structured matrices.
III Parallel Selected Inversion
This section introduces an algorithm for inverting symmetric matrices, with a particular emphasis on arrowhead structures, using a tile-based approach grounded in the Cholesky decomposition $A=LL^{T}$. This Cholesky-based methodology consists of three main phases: heuristic ordering to optimize the preservation of sparsity, symbolic factorization to determine the structure and dependencies within the factorization, and numerical factorization to compute the actual values of the matrix elements. Further details on the Cholesky-based implementation are provided in the context of sTiles [14].
Figure 2: Illustration of Cholesky factorization and inversion patterns of selected tiles for dense and arrowhead matrices
Building on this Cholesky-based framework, the matrix inversion process is structured as follows:
  1. 1.
    The selection step begins with the user specifying a list of matrix elements, identified by their indices $(i,j)$, for which the inverse is required. These indices are mapped to their corresponding tiles using the same compressed tile format employed during the Cholesky decomposition. This step ensures consistency in data representation and prepares the structure for subsequent computations.
  2. 2.
    A symbolic inversion step follows, during which all necessary dependencies for the selected elements are identified and incorporated. This step guarantees the completeness and accuracy of the subsequent computations.
  3. 3.
    Finally, the numerical inversion is performed, computing the desired inverse elements. This step leverages the tile-based structure to maximize computational efficiency.
Next, we delve into the implications of this process by examining different matrix structures and their corresponding inversion patterns, with a particular focus on dense and arrowhead matrices.
III-A Selected Elements of the Inverse
Consider the task of computing the inverse for a selected set of element pairs within a matrix. The matrix is partitioned into tiles, and as an initial step, these elements are mapped to their corresponding tiles, referred to as selected tiles. Although this approach necessitates computing the inverse for all elements within a tile, even if only a subset is of primary interest, it provides significant advantages at negligible cost in memory and computational complexity. This computation is essential due to fill-in, as many elements within these tiles become populated during the inversion process. Additionally, from a memory efficiency standpoint, using tiles improves cache utilization, thereby enhancing overall computational performance.
Our focus is on computing the selected inverse of structured matrices, with particular attention to arrowhead patterns. Figure 2 presents ten illustrative cases divided into two groups: fully dense matrices (cases 1-5) and arrowhead matrices (cases 6-10). For each case, we show the original matrix $A$, its Cholesky factor, the selected tiles requested for inversion, and the resulting inversion pattern.
In cases 1-3, the original matrix is fully dense, and the selected tiles include the diagonal. As a result, the inversion covers the entire lower triangle, producing a full inverse. This behavior reflects the high computational cost of inverting dense matrices when the diagonal is involved. Case 6 mirrors this behavior in the arrowhead setting. Although the structure is sparse, selecting the entire matrix leads to full inversion, analogous to case 1.
Cases 7 and 8 represent arrowhead matrices where the selected tiles form an arrowhead structure that includes the diagonal. In these cases, the resulting inverse retains the same arrowhead structure. This behavior closely matches that of cases 2 and 3 in the dense setting, where the selected tiles include the diagonal and do not extend beyond the original non-zero pattern.
Cases 4-5 and 9-10 show a consistent pattern across both dense and arrowhead matrices: the selected tiles do not include any diagonal tiles. Consequently, only a minimal subset of the inverse is computed, and the cost remains low. These cases highlight how omitting the diagonal and restricting selection to isolated tiles significantly reduces the computational effort required for inversion.
Our focus is on the arrowhead matrix in which the selected pattern matches the Cholesky pattern, specifically case 7. This formulation can be adapted to case 6 when needed. We next present the Directed Acyclic Graph (DAG) representing the inversion process for cases 2 and 7.
III-B Directed Acyclic Graph
Figure 3: Directed acyclic graph (DAG) representations for selected inversion on matrices of size $6\times 6$ tiles. The upper plot corresponds to Case 2 (dense matrix) in Figure 2, while the lower plot corresponds to Case 7 (arrowhead matrix). These DAGs visualize task dependencies and parallelism within each structure.
The Directed Acyclic Graph (DAG) captures dependencies between computations, enabling efficient parallel execution and optimized resource allocation. By organizing tasks with well-defined precedence, the DAG ensures that independent computations can proceed concurrently. The DAG underlying our approach is constructed using four key computational kernels—TRSM, LAUUM, GEMM, and TRMM—each applied to specific tiles $\Sigma_{i,j}$, corresponding to the tile at position $(i,j)$. The definitions and roles of these operations are described in detail in Section II.
As illustrated in Figure 3, the DAGs for full and arrowhead matrix inversions (Cases 2 and 7, respectively) exhibit notable differences in structure and parallelism. The DAG for full matrix inversion is characterized by a large width, indicating many tasks that could, in principle, be executed concurrently. In the case of the arrowhead structure, the number of nodes is significantly reduced due to the pruning of unnecessary computations when only selected tiles are targeted for inversion. This leads to a more compact DAG with a level of concurrency depending on which selected tile inversions are needed and reduced overall computational workload. Importantly, the critical path length remains the same in both cases, consisting of six sequential operations along the longest dependency chain for the studied structured matrices.
III-C Two-Phase Algorithm for Selected Inversion
To parallelize Algorithm 1, we adopt a two-phase approach designed to minimize interdependencies between cores, thereby reducing idle times and enhancing parallel efficiency. Beyond parallelism, this approach also prunes unnecessary computations by identifying and skipping tiles that do not contribute to the user-specified subset of the inverse. By leveraging the structure of the selected tiles, we avoid redundant operations while preserving correctness. The division of the algorithm into distinct phases enables better task organization and facilitates scalable execution. A critical aspect of this strategy is the use of static load balancing, where tasks are preassigned to cores during preprocessing. This ensures an even distribution of the workload and significantly reduces runtime scheduling overhead.
Phase 1: Independent Tile Computations
In this phase, each core operates independently on specific columns of the matrix, determined by its thread ID and the total number of participating cores. The columns are processed in a round-robin manner, with the column index assigned based on the core’s thread ID modulo the total number of cores. This approach ensures a balanced and evenly distributed workload across all cores. During this phase, the values in the upper triangular matrix $L^{T}$ are updated to prepare for the inversion in the second phase. Additionally, the intermediate results of the selected inversion are progressively stored in the matrix $\Sigma$, which will eventually contain the selected inverse.
To define the structure, we introduce the notation $j\in\text{neighbors}(i)$ if and only if tile $X_{ij}^{T}$ or tile $X_{ji}^{T}$ ($X=L$ or $X=\Sigma$ ) corresponds to a non-zero tile. A non-zero tile may initially be zero but is added during the symbolic inversion phase and is fully updated by the end of the computation. The algorithm for this phase is detailed in Algorithm 2.
Algorithm 2 Parallel selected inversion - phase 1
Given: Matrix $A$ is partitioned into $N\times N$ tiles and expressed as $A=LL^{T}$, where $L^{T}$ is the upper triangular factor.
Note: $\Sigma$ is the matrix where the selected inverse is stored.
Initialization: $i\leftarrow N-1-\texttt{thread ID}$
while $i\geq 0$ do
     for all $j\in\text{neighbors}(i)$ and $i\leq j<N-1$ do
         if $i==j$ then
              $\Sigma_{ii}^{T}\leftarrow\texttt{TRSM}(L_{ii}^{T},I)$
         else
              $L_{ij}^{T}\leftarrow\texttt{TRMM}(L_{ii}^{T},L_{ij}^{T})$
         end if
     end for-
[…]
end while
Phase 2: Dependent Tile Computations
Once all cores have completed their assigned tasks in phase 1, the algorithm transitions to phase 2, where the intermediate results in $\Sigma$ are further processed to compute the final selected inverse.
In phase 2, the intermediate matrix $\Sigma$ is updated using a parallel, task-based approach that respects data dependencies across cores. Each core operates on a specific column $i$, determined by its thread ID, and iterates through dependent rows $j$ and neighboring blocks $k$. Diagonal blocks are updated using contributions from their neighbors, while off-diagonal blocks are computed based on interdependencies between neighboring tiles. The execution follows an asynchronous model, where each core advances independently based on the readiness of required data. A progress tracking mechanism, core_progress, provides lightweight, fine-grained synchronization to enforce correct task ordering without requiring global barriers. This design enables efficient parallelization while handling the complex data dependencies inherent to the selected inversion process.
Algorithm 3 Parallel selected inversion - phase 2
Initialization: $i\leftarrow N-1-\texttt{thread ID}\texttt{ and }\texttt{set = false;}$;
while $i\geq 0$ do
     for $j=N-1$ to $i$ step $-1$ do
         if $i==j$ then
              $\Sigma_{ii}\leftarrow\texttt{LAUUM}(\Sigma_{ii})$
              for all $k\in\text{neighbors}(i)$ and $i<k<N-1$ do
                  $\Sigma_{ii}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{ik}^{T},\Sigma_{ii})$
              end for
              Set core_progress[i,i] = 1
         else
              if $(i+1)\leq(N-1)$ then
                  set = true; ii = 0; jj = 0;
              end if
              for all $k\in\text{neighbors}(j)$ and $i<k<N-1$ do
                  if $i\in\text{neighbors}(j)\cap j\in\text{neighbors}(k)$ and $i<j$ then
                       if $k>j$ then
                           while core_progress[j,k] $\neq$ 1 do
                                /* Wait */
                           end while
                           $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{kj},\Sigma_{ij})$
                           ii = i; jj = j;
                       else
                           while core_progress[k,j] $\neq$ 1 do
                                /* Wait */
                           end while
                           $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{jk}^{T},\Sigma_{ij})$
                           ii = i; jj = j;
                       end if
                  end if
              end for
              if set then
                  Set core_progress[ii,jj] = 2;
              end if
         end if
     end for
[…]
end while
Figure 4: Directed acyclic graphs for full and arrowhead matrix inversions using 2 cores and 4 cores, with each color representing tasks for each core.
The Directed Acyclic Graphs (DAGs) in Figure 4 directly represent the task distribution and parallelism strategies implemented in the two-phase algorithm detailed above. The graphs demonstrate how tasks are assigned to different cores, with each color representing the tasks handled by a specific core. While this visualization uses matrices of size $6\times 6$ tiles, the distribution and parallelism would become even more apparent for larger matrices, where the increased number of tasks leads to a more complex workload distribution.
III-D GPU Acceleration for Parallel Selected Inversion
The GPU implementation of parallel selected inversion in sTiles follows a similar tile-based approach as its CPU counterpart, with adaptations to leverage the massive parallelism and high throughput of modern GPUs. Given that selected inversion focuses only on specific elements of the inverse, we assume that the selected tiles can fit entirely within a single GPU’s memory, ensuring efficient execution without the need for frequent data transfers between the CPU host and the GPU device.
The same tile size criteria used in the GPU implementation of Cholesky factorization in sTiles is adopted for the GPU version of the selected inversion to maintain consistency in memory access patterns and workload distribution. Each tile is mapped to a dedicated CUDA stream, enabling concurrent execution of independent tasks across multiple GPU compute units.
To optimize performance and minimize data transfer overhead, the entire matrix, along with its factorization, is fully copied to GPU memory before any computations begin. All computational kernels in the CPU implementation, such as Cholesky factorization, triangular solves, symmetric rank-k updates, and matrix multiplications, are replaced with their respective cuBLAS and cuSOLVER implementations. The key operations include:
  • Triangular solve: cublasDtrsm
  • Triangular matrix multiply: cublasDtrmm
  • Matrix multiplication: cublasDgemm
Once all computations are completed, the results are transferred back to the CPU for further processing or storage. Since data movement between CPU and GPU is a major bottleneck, the design ensures that all required computations are performed on the GPU before any data is transferred back, maximizing throughput and reducing unnecessary communication overhead.
This GPU-accelerated implementation of parallel selected inversion in sTiles significantly improves performance for structured matrices by fully utilizing GPU resources, minimizing data transfers, and leveraging parallel execution through CUDA streams.
While our current implementation assumes that the selected tiles fit in GPU memory, the algorithm could be extended to support out-of-core execution for larger matrices that exceed GPU capacity. This would involve overlapping data movement and computation while maintaining efficient static scheduling. Such strategies have been successfully applied in the context of out-of-core Cholesky factorization on modern GPU architectures, as demonstrated in [17]. Incorporating similar techniques into sTiles would enable scalable selected inversion on next-generation heterogeneous systems such as the NVIDIA Grace Hopper Superchip with a unified memory subsystem on the CPU host.
IV Performance Evaluation and Experimental Results
In this section, we present a comprehensive evaluation of our GPU-accelerated parallel selected inversion implementation using sTiles. While we compare its CPU implementation against an existing state-of-the-art CPU-only approach, our goal for the GPU implementation is to demonstrate the impact of the sTiles fine-grained algorithmic approach on massively parallel hardware accelerators.
IV-A Experimental Setup and Software
Our computational experiments were conducted on two high-performance computing (HPC) systems, each designed to assess the performance of sTiles under different computational paradigms:
  • CPU Server: A dual-socket 26-core system featuring Intel® Xeon® Gold 6230R processors, each equipped with 52 cores total operating at 2.10 GHz, with a total L3 cache of 71.5 MB.
  • GPU Node: A dedicated compute node incorporating 64 AMD EPYC 7713 CPU cores running at 1.99 GHz, complemented by a single NVIDIA A100-SXM4 GPU, which operates at 1.16 GHz and includes 80 GB of high-bandwidth memory (HBM2).
This setup facilitates a comprehensive evaluation of sTiles across CPU-only and GPU-accelerated configurations, providing insights into its computational efficiency and scalability.
To benchmark our parallel selected inversion algorithm, we compare it against Panua-PARDISO 8.2, a state-of-the-art direct solver for sparse linear systems utilizing $LL^{T}$ factorization, optimized for shared-memory parallelism. This solver employs advanced numerical techniques to ensure efficient factorization and inversion, making it a robust reference for our comparative analysis.
Additionally, SelInv 2.0.0 was initially considered for inclusion in our performance assessment. However, several issues precluded a fair and meaningful comparison: (1) the software is no longer actively maintained, (2) the only functional configuration we identified runs with a single MPI process and mandates a specific ordering option, and (3) certain ordering configurations fail to work. Moreover, SelInv 2.0.0 relies on SuperLU for its inverse computation backend, whereas our approach assumes $LL^{T}$ factorization. Due to these constraints, we excluded SelInv 2.0.0 from our evaluation.
IV-B Performance Evaluation
TABLE I: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 1. These matrices reflect the arrowhead structures that commonly arise in INLA-based models [4].
ID Size Bandwidth Arrowhead Thickness Density (%)
1 10,010 100 10 0.408
2 10,010 200 10 0.605
3 10,010 300 10 0.643
4 10,200 100 200 3.938
5 10,200 200 200 4.032
6 10,200 300 200 4.066
7 100,010 1000 10 0.121
8 100,010 2000 10 0.219
9 100,010 3000 10 0.258
10 100,200 1000 200 0.498
11 100,200 2000 200 0.597
12 100,200 3000 200 0.637
13 500,010 1000 10 0.024
14 500,010 2000 10 0.044
15 500,010 3000 10 0.052
16 500,200 1000 200 0.100
17 500,200 2000 200 0.120
18 500,200 3000 200 0.128
Table I summarizes the structured matrices utilized in our evaluation, encompassing variations in size, bandwidth, and sparsity levels. These matrices are carefully chosen to represent practical applications, such as statistical models with structured sparsity patterns. Specifically, the arrowhead thickness corresponds to the number of fixed effects in statistical models, with values ranging from 10 (moderate case) to 200 (extreme case). The benchmarking aims to assess the computational efficiency of sTiles in handling these structures and to compare its performance against Panua-PARDISO 8.2, a state-of-the-art solver.
Our benchmarking methodology consists of the following steps:
  1. 1.
    Factorization and Selected Inversion: Cholesky factorization is performed, followed by selected inversion using sTiles. The execution time is compared against Panua-PARDISO 8.2.
  2. 2.
    Performance Evaluation: Execution times are recorded across increasing core counts (1, 2, 4, 8, 16, 32, and 52 cores, the maximum available on our test system). The best execution time for each configuration is selected.
  3. 3.
    Performance Visualization: Results are analyzed in terms of absolute runtime and relative speedup compared to Panua-PARDISO.
Figure 5: Performance comparison of sTiles and Panua-PARDISO across different matrix configurations. Each subfigure represents a different matrix size category: small (top), medium (middle), and large (bottom).
Figure 5 demonstrates the performance advantage of sTiles over Panua-PARDISO across different matrix sizes. The performance gap becomes more pronounced as matrix size increases (IDs 7-18), with sTiles leveraging dense tiles to mitigate computational overhead and memory bandwidth limitations. In some cases, sTiles achieves up to 13.49× speedup compared to Panua-PARDISO.
IV-C Scalability Evaluation
Figure 6: Execution times of sTiles and PARDISO for matrices of size $10K$ (left) and $500K$ (right) across different core counts. The y-axis is in logarithmic scale.
Figure 6 presents the execution times of sTiles and Panua-PARDISO across different core counts for two representative matrix sizes: $10K$ and $500K$. For the smaller matrix, both solvers exhibit strong scalability, maintaining comparable performance across all configurations. However, for the larger matrix, sTiles demonstrates superior scalability, achieving consistently lower execution times as the core count increases. This performance advantage is attributed to sTiles’ efficient exploitation of parallelism.
IV-D Impact of Sparsity on Performance
The impact of sparsity on the computational performance of selected inversion is analyzed using a set of structured matrices with varying densities. Table II summarizes the properties of these matrices, including size, bandwidth, arrowhead thickness, and density. The density values, which exclude the arrowhead portion, range from 0.010% to 4.101%, providing a broad spectrum of sparsity levels for evaluation.
Figure 7 presents the inverse computation times for both PARDISO and sTiles solvers across matrices with increasing density. The results illustrate a clear distinction in solver behavior depending on the sparsity characteristics. For matrices with very low density (below 0.1%), PARDISO exhibits a faster performance than sTiles, primarily due to its optimized multifrontal structure for handling highly sparse matrices. However, as the density increases beyond this threshold, the performance of sTiles stabilizes, while PARDISO experiences significant computational overhead.
The results indicate that sTiles maintains consistent computational times across a wide range of density values, demonstrating its robustness for handling moderately sparse to dense matrices. In contrast, PARDISO’s runtime increases substantially with density, reflecting the growing complexity of fill-in and symbolic factorization in direct solvers. For high-density matrices (greater than 1%), sTiles outperforms PARDISO, making it a more scalable approach for dense and structured sparse matrices.
This analysis highlights the advantage of the sTiles approach in scenarios where matrix density increases while preserving a structured sparsity pattern. The ability of sTiles to sustain lower computational cost across different density regimes makes it an attractive alternative for large-scale scientific computing applications where memory and time efficiency are critical constraints.
TABLE II: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 2. Density values exclude the arrowhead part.
Matrix Size = 10,004   Arrowhead Thickness = 4
Bandwidth = 1500 Bandwidth = 3000
ID Density (%)    ID Density (%)
19 0.010    34 0.010
20 0.018    35 0.026
21 0.031    36 0.051
22 0.054    37 0.076
23 0.095    38 0.092
24 0.139    39 0.255
25 0.181    40 0.339
26 0.227    41 0.417
27 0.266    42 0.501
28 0.309    43 0.584
29 0.354    44 0.668
30 0.398    45 0.749
31 0.437    46 0.828
32 0.871    47 1.651
33 2.153    48 4.101
Figure 7: Selected inverse computation times of sTiles and PARDISO across varying matrix densities. The top plot illustrates results for matrices with a bandwidth of 1500, while the bottom plot corresponds to matrices with a bandwidth of 3000. The x-axis represents matrix IDs, with density values displayed beneath each ID, while the y-axis uses a logarithmic scale to capture the variations in computation time.
IV-E Full Matrix Inversion Performance
This section examines the performance of sTiles in full matrix inversion and compares it against PLASMA [16, 15], highlighting the impact of scheduling strategies. PLASMA employs dynamic scheduling, while sTiles uses static scheduling to increase data locality and minimize scheduling overhead.
Figures 8 and 9 present the execution times of both libraries across different tile sizes and core counts for matrix ID 5 from Table I. The main observations are:
  • sTiles consistently outperforms PLASMA, achieving lower execution times across all configurations.
  • PLASMA suffers from performance degradation beyond 16 cores, particularly for small tile sizes (40 and 80). This is attributed to dynamic scheduling overhead and thread contention.
  • sTiles scales efficiently, maintaining a steady decrease in execution time up to 52 cores, demonstrating better parallel efficiency.
  • Tile size sensitivity is lower in sTiles, meaning it achieves strong performance across different tile sizes without requiring extensive tuning.
Figure 8: Execution times of PLASMA for full matrix inversion across different tile sizes and core counts for matrix ID 5.
Figure 9: Execution times of sTiles for full matrix inversion across different tile sizes and core counts for matrix ID 5.
These results indicate that sTiles remains efficient even in full matrix inversion scenarios, maintaining strong scalability and robust performance across all tile sizes, while PLASMA’s dynamic scheduling leads to a large time complexity at high core counts. To further emphasize this, we plot the full inversion comparison using the best tile size for each library (PLASMA: 280, sTiles: 320) in Figure 10.
Figure 10: Full inversion comparison using the best tile size for each library (PLASMA: 280, sTiles: 320). sTiles consistently achieves lower execution times across all core counts.
IV-F Accelerating Selected Inversion on GPU
To evaluate the benefits of GPU acceleration, we performed selected inversion on a large matrix of size 50,010, with a bandwidth of 15,000, an arrowhead thickness of 10, and a density of 0.3123. This configuration represents a computationally intensive problem well-suited for GPU execution, as the large bandwidth leads to dense blocks and a high ratio of floating-point operations to memory accesses.
Following the experimental setup in [14], we used a tile size of 120 for CPU computations and a larger tile size of 600 for GPU computations. While the GPU hardware itself offers inherent performance advantages, additional gains are achieved through preprocessing optimizations during factorization. Notably, differences in tile size influence the structure of the tiled matrix, symbolic factorization complexity, and the distribution of computational tasks across resources.
TABLE III: CPU vs GPU performance comparison for the two-phase selected inversion.
Cores/Streams CPU Time (s) GPU Time (s)
1 533.062 7.044
2 282.714 4.524
4 142.883 4.366
8 75.225 3.954
16 39.575 4.111
32 26.861 3.562
64 20.382 3.949
All GPU results were obtained using the NVIDIA A100 GPU available on the GPU node described in Section IV-A, while CPU results were measured on the dual-socket 26-core Intel Xeon server.
As shown in Table III, the GPU achieved a runtime of 3.949 seconds, compared to 20.382 seconds on 64 CPU cores, yielding a speedup of 5.12$\times$.
These results highlight the efficiency of GPU acceleration for large-scale structured matrix problems, particularly when combined with tile-aware design choices that exploit the high arithmetic intensity and parallelism available on modern GPU architectures.
The observed GPU speedup aligns well with expectations derived from the Roofline Performance Model, which suggests a theoretical peak performance ratio of up to 20$\times$ between CPU and GPU (1 TFLOP/s vs. 20 TFLOP/s). In practice, however, the speedup is moderated by host-to-device data transfer overheads and the inherent limitations of workload concurrency. While the GPU architecture is highly effective for throughput-oriented parallel computations, its ability to reach peak performance depends on the volume and granularity of the workload. When concurrency is limited or data movement becomes a bottleneck, the achievable speedup is reduced, although it may still indicate performance superiority.
V Conclusion
In this work, we introduced an efficient GPU-accelerated parallel selected inversion algorithm for structured matrices using sTiles. Our approach leverages a tile-based methodology that efficiently handles structured sparsity, particularly for arrowhead matrices, ensuring both computational efficiency and parallel scalability. By adopting a two-phase algorithm, we minimized interdependencies between computational tasks, allowing for efficient static scheduling and improved parallel execution across both CPU and GPU architectures.
The implications of this work extend beyond selected inversion, as the tile-based framework introduced in sTiles can be extended to other numerical linear algebra problems that benefit from hybrid sparse-dense computations, offering a wide range of functionalities. Future work includes extending the framework to enabling multi-GPU support, supporting distributed-memory architectures, and optimizing for even larger-scale problems encountered in scientific computing, Bayesian inference, and high-dimensional statistical modeling. By continuously refining the sTiles framework, we aim to push the computational boundaries of structured matrix computations and enhance its applicability in real-world high-performance computing applications.